- Press Release
- Oct 6, 2022
NOAA GOES SATOPS Morning Report 27 October 2000
SATOPS MORNING REPORT October 27, 2000
No change in the status of the POES spacecraft.
NOAA-16 operations nominal.
The product quality problem with data from WCDAS was resolved at 1945z on
Thursday, October 26, by cycling the Polar Frame Synch (PFS3). After
troubleshooting ground system components, PFS3 was identified at the source of
noise on the Wallops output line (DPSS Interface 1) from SOCC to CEMSCS.
Recycling power on PFS3 cured the problem, and CEMSCS is now receiving good data
A NOAA-15 solar array adjustment is planned for Monday to further shade the
IMP, followed by a few passes where the MIRP and AVHRR will be synch’ed while
within sight of a ground station (~10 minute durations) to collect synch delta
(timing) data. During the week, the AVHRR will be cycled between high and low
Overnight, the NOAA-15 HIRS was switched into high power mode via on-board
MACRO commanding due to the filter wheel current threshold being exceeded.
Engineering investigations will continue today.
NOAA-15 rev 12755 / W at 1220z on October 26: G2B lost (bit synch lock but
not frame synch). Data not recoverable on subsequent revs due to recorder
NOAA-14 rev 30018 / F at 1927z on October 26: G3A noisy due to AGC fade. 35%
of the data was captured. Data not recoverable on subsequent revs due to
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
On F15/OPS49 Rev 4519, SCP 1 dropped into BASIC. SCP 1 will be commanded back
into SINGLE after both SCP’s converge.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
At 2350z on Thursday, October 26, and at 0003z and 0023z on Friday, October
27, GOES-8 frame breaks were observed and attributed to high BER at CDA (unknown
Missed the GOES-10 2130z Southern Hemi image due to a T1 outage.
GOES-10 East/West maneuver # 14 was executed nominally this morning at 0924z.
The commanded thruster firing duration was 8.3 seconds.
GOES-9 and GOES-11 remain in the Z-Axis Precession (ZAP) storage mode at ~105
GOES operations were nominal over the past 24 hours.
Questions or comments to firstname.lastname@example.org